OpenStudy (anonymous):

***Help for a medal and fan!*** Write the following expression as a function of a positive acute angle. cos 125° a)cos55° b)-cos55° c)cos(-55)°

OpenStudy (anonymous):

\[\cos(180-\theta)=-\cos(\theta)\]

OpenStudy (anonymous):

cos(180-125)=-cos(55)

OpenStudy (anonymous):

so in my calculator I just type in cos(180-125) then will that give me my answer? can you explain please? @nitz

OpenStudy (anonymous):

u dont need a calculator

OpenStudy (anonymous):

The acute angle is the small angle which is less than 90°.

OpenStudy (anonymous):

ok... so i just subtract ? and how will i know if the cos is negative or not

OpenStudy (anonymous):

to convert the expression........ \[\cos(180-\theta)=-\cos(\theta)\]

OpenStudy (anonymous):

\[\cos(\theta) \] is negaTIVE IN SECOND QUADRANT....

OpenStudy (anonymous):

and \[(180-\theta)\] is the angle in second quadrant

OpenStudy (anonymous):

okay i get it.... thanks for your help!!!

OpenStudy (anonymous):

<90 first quadrant \[90<\theta<180\] 2nd quadrant \[180<\theta<270\] third quadrant \[270<\theta<360 \] fourth quadrant

OpenStudy (anonymous):

cos is negative in second and third quadrant....

OpenStudy (anonymous):

thanks! :)